Tangshan Antiques has a beta of 1.40, the annual risk-free rate of interest is currently 10 percent, and the required return on the market portfolio is 16 percent. The firm estimates that its future dividends will continue to increase at an annual compound rate consistent with that experienced over the 2016-2019 period.

Year 2016 Dividend $2.70 2017 2.95 2018 3.25 3.40 2019

Estimate the value of Tangshan Antiques stock at the end of 2019. After the 2019 dividend has been paid. Round your final answer to two decimal places.
Answer #1

first let us know the growth rate of dividend.

2016 year dividend *(1+r)^n = 2019 dividend.

here,

r is to be found out

n = 3 years.

=>2.70 *(1+r)^3 = 3.40

=>(1+r)^3 =>1.259259

=>1+r = (1.259259)^(1/3)

=>1+r=1.0798706

=>r=0.0798706

=>r=8.00%...(rounded off).

so growth rate =8%.

now,

let us know the required rate of return = risk free rate + beta*(market return - risk free rate)

=>10%+ 1.40*(16-10)

=>18.40%.

price of the stock = 2019 dividend *(1+ growth rate) / (required rate - growth rate)

=>3.40*(1.08) / (0.1840-0.08)

=>$35.31.